OpenStudy (kaylak12345):

solve the system by elimination -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (kaylak12345):

@tigerlover

OpenStudy (kaylak12345):

i need help

OpenStudy (jango_in_dtown):

@kaylak12345 Gaussian elimination?

OpenStudy (kaylak12345):

yes

OpenStudy (jango_in_dtown):

Write the augmented matrix first.

OpenStudy (kaylak12345):

-2 2 3 0 -2 -1 1 -3 2 3 3 5

OpenStudy (jango_in_dtown):

apply R2=R2-R1 and R3=R3+R1. then the result is?

OpenStudy (kaylak12345):

what is r

OpenStudy (jango_in_dtown):

row

OpenStudy (jango_in_dtown):

2nd row(new)=2nd row(old)-1st row(old). 3rd row(new)= 3rd row(old)+1st row(old)

OpenStudy (kaylak12345):

2 -1 -2 -3 for r2?

OpenStudy (jango_in_dtown):

try again..

OpenStudy (jango_in_dtown):

what i am trying is to make all the elements of the first column expect the first to be zero

OpenStudy (kaylak12345):

now m lost

OpenStudy (kaylak12345):

could you example with row 1 and 2 andlet me try 2 and 3

OpenStudy (kaylak12345):

0 -1 -2 -3 im lost

OpenStudy (kaylak12345):

i have to show my work

OpenStudy (jango_in_dtown):

check the attachment

OpenStudy (jango_in_dtown):

I will help you..

OpenStudy (kaylak12345):

so what after that

OpenStudy (jango_in_dtown):

did you understood how i got it?

OpenStudy (kaylak12345):

somewhat

OpenStudy (jango_in_dtown):

Its what I told, we want to make all the elements below the first element of column 1 to be zero

OpenStudy (jango_in_dtown):

thats why i told you to subtract the 1st row from 2nd and add the 3rd and first row elements

OpenStudy (kaylak12345):

oh now im getting it still kinda lost but understanding it btter

OpenStudy (jango_in_dtown):

okay... the basic concept is to make an upper triangular matrix and then use back substitution

OpenStudy (jango_in_dtown):

now we will make the element below -3 of second row 0. i.e. we want to make the 5 in the second column to be 0

OpenStudy (kaylak12345):

5-3=2-2=0

OpenStudy (kaylak12345):

right

OpenStudy (jango_in_dtown):

for it. apply R3=R3+(5/3)R2

OpenStudy (jango_in_dtown):

the 3rd row will be 0 0 8/3 0

OpenStudy (kaylak12345):

was i right

OpenStudy (jango_in_dtown):

no 5-3 is not 2-2

OpenStudy (kaylak12345):

−2+2y+3z=0 2y+3z−2+2=0+2 2y+3z=2 2y+3z+−3z=2+−3z 2y=−3z+2 2y/2 =−3z+2/2 y =−3/2z + 1

OpenStudy (kaylak12345):

is this y

OpenStudy (kaylak12345):

−2x−y+z=−3 −2x−y+z+y=−3+y −2x+z=y−3 −2x+z+−z=y−3+−z −2x=y−z−3 −2x/−2 = y−z−3/−2 x =−1/2y +1/2z +3/2

OpenStudy (kaylak12345):

and x

OpenStudy (jango_in_dtown):

elisaab gave the solution in the attachment